OpenStudy (anonymous):

Suppose x>8, which of the following must be true? (Ther's a Drawing)

OpenStudy (anonymous):



jhonyy9 (jhonyy9):

3x-1=2x+7 3x-2x=7+1 x=8 so if x>8 so than AC not can being equal EF or AB not can being equal ED hope so much that this will help